# Key-Point-Driven Data Synthesis with its Enhancement on Mathematical Reasoning

### v2

@@ Abstract
Large language models (LLMs) have shown great potential in complex reasoning tasks, yet their performance is often hampered by the scarcity of high-quality and reasoning-focused training datasets. Addressing this challenge, we propose Key-Point-Driven Data Synthesis (KPDDS), a novel data synthesis framework that synthesizes question-answer pairs by leveraging key points and exemplar practices from authentic data sources. KPDDS ensures the generation of novel questions with rigorous quality control and substantial scalability. As a result, we present KPMath, an extensive synthetic dataset tailored for mathematical reasoning, comprising over 800K question-answer pairs. Utilizing KPMath and augmenting it with additional reasoning-intensive corpora, we create the comprehensive KPMath-Plus dataset. The fine-tuned DeepSeekMath model on KPMath-Plus achieves zero-shot PASS@1 accuracies of 83.9% on GSM8K and 48.8% on MATH, and also reaches promising performance on other math reasoning datasets, outperforming competitors in the 7B to 70B range.
@@ 1 Introduction
Utilizing the training sets of the MATH (Hendrycks et al., 2021b) and GSM8K (Cobbe et al., 2021b) datasets as foundational data, we developed a novel dataset named KPMath. Our training corpus was further enriched by integrating a series of mathematical reasoning datasets, leading to the creation of a comprehensive training dataset, KPMath-Plus. By fine-tuning the Mistral-7B model (Jiang et al., 2023) on KPMath-Plus, we achieved zero-shot PASS@1 accuracies of 82.1% on the GSM8K test set and 46.8% on the MATH test set, culminating in a promising average of 75.4% across six math reasoning datasets. This performance exceeds that of all competitors within the 7B to 70B model size range, thus achieving state-of-the-art status. In the Hungarian Exam Score test, the KPMath-Plus-Mistral-7B model also outperforms the majority of models, indicating its competitive performance.
@@ 4.4 Baselines
We present results from a range of state-of-the-art (SoTA) proprietary LLMs, including OpenAI’s GPT-4 (OpenAI, 2023), ChatGPT (gpt-3.5-turbo), Google’s PaLM-2(Anil et al., 2023), and Anthropic’s Claude-2(Anthropic, 2023). Regarding open-source models, we consider base models such as LLaMA-2(Touvron et al., 2023), DeepSeekMath(Shao et al., 2024), Mistral(Jiang et al., 2023), and Llemma (Azerbayev et al., 2023). Supervised Fine-Tuning (SFT) employs CoT rationales from the original GSM8k and MATH dataset (15k samples) for fine-tuning. We also showcase the performance of advanced models using SFT or RLHF on various mathematical reasoning datasets, including MAmmoTH (Yue et al., 2024), WizardMath (Luo et al., 2023a), Platypus-2 (Lee et al., 2023), MetaMath (Yu et al., 2023b) and MMIQC (Liu & Yao, 2024).
@@ 4.5 Main Results
Table 1 presents the results on six widely-used mathematical benchmarks, highlighting several key observations: KPMath-Plus significantly enhances the performance of multiple base models, with average accuracy improvements ranging from 10.6% to 36.9%. The KPMath-Plus-DeepSeekMath model achieves zero-shot PASS@1 accuracies of 83.9% on GSM8K and 48.8% on MATH, and also reaches promising performance on other math reasoning datasets, outperforming competitors in the 7B to 70B range.

### v3

@@ Abstract
Large language models (LLMs) have shown great potential in complex reasoning tasks, yet their performance is often hampered by the scarcity of high-quality and reasoning-focused training datasets. Addressing this challenge, we propose Key-Point-Driven Data Synthesis (KPDDS), a novel data synthesis framework that synthesizes question-answer pairs by leveraging key points and exemplar practices from authentic data sources. KPDDS ensures the generation of novel questions with rigorous quality control and substantial scalability. As a result, we present KPMath, an extensive synthetic dataset tailored for mathematical reasoning, comprising over 800K question-answer pairs. Utilizing KPMath and augmenting it with additional reasoning-intensive corpora, we create the comprehensive KPMath-Plus dataset. The Qwen1.5-72B model, fine-tuned on KPMath-Plus, achieves 87.0% PASS@1 accuracy on GSM8K and 58.3% on MATH, surpassing competitors in the 7B to 70B range and best commercial models like GPT-4 across multiple math reasoning datasets.
@@ 1 Introduction
Utilizing the training sets of the MATH (Hendrycks et al., 2021b) and GSM8K (Cobbe et al., 2021b) datasets as foundational data, we developed a novel dataset named KPMath. Our training corpus was further enriched by integrating a series of mathematical reasoning datasets, leading to the creation of a comprehensive training dataset, KPMath-Plus. By fine-tuning the Qwen1.5-72 model (Bai et al., 2023) on KPMath-Plus, we achieved zero-shot PASS@1 accuracies of 87.0% on the GSM8K test set and 58.3% on the MATH test set, culminating in a promising average of 81.5% across six math reasoning datasets. This performance exceeds that of all competitors within the 7B to 70B model size range and best commercial models like GPT-4. In the Hungarian Exam Score test, the KPMath-Plus-Mistral-7B model also outperforms the majority of models, indicating its competitive performance.
@@ 4.2 Implementation Details
| KPMath-Plus | Llama-2 | 70B | ✓ | 87.4 | 48.6 | 81.2 | 75.1 | 89.0 | 95.4 | 79.4 (+21.2) |
| Qwen1.5 | - | 72B | ✗ | 77.6 | 38.2 | 82.5 | 52.0 | 85.1 | 95.9 | 71.9 |
| KPMath-Plus | Qwen1.5 | 72B | ✓ | 87.0 | 58.3 | 82.1 | 76.7 | 89.2 | 95.5 | 81.5 (+9.6) |
@@ 4.4 Baselines
We present results from a range of state-of-the-art (SoTA) proprietary LLMs, including OpenAI’s GPT-4 (OpenAI, 2023), ChatGPT (gpt-3.5-turbo), Google’s PaLM-2(Anil et al., 2023), and Anthropic’s Claude-2(Anthropic, 2023). Regarding open-source models, we consider base models such as LLaMA-2(Touvron et al., 2023), DeepSeekMath(Shao et al., 2024), Mistral(Jiang et al., 2023), Llemma (Azerbayev et al., 2023), and Qwen1.5(Bai et al., 2023). Supervised Fine-Tuning (SFT) employs CoT rationales from the original GSM8k and MATH dataset (15k samples) for fine-tuning. We also showcase the performance of advanced models using SFT or RLHF on various mathematical reasoning datasets, including MAmmoTH (Yue et al., 2024), WizardMath (Luo et al., 2023a), Platypus-2 (Lee et al., 2023), MetaMath (Yu et al., 2023b) and MMIQC (Liu & Yao, 2024).
@@ 4.5 Main Results
Table 1 presents the results on six widely-used mathematical benchmarks, highlighting several key observations: KPMath-Plus significantly enhances the performance of multiple base models, with average accuracy improvements ranging from 10.6% to 36.9%. The KPMath-Plus-Qwen1.5-72B model achieves zero-shot PASS@1 accuracies of 87.0% on GSM8K and 58.3% on MATH, and also reaches promising performance on other math reasoning datasets, outperforming competitors in the 7B to 70B range.